### v1

# Promise and Peril: Stellar Contamination and Strict Limits on the Atmosphere Composition of TRAPPIST-1c from JWST NIRISS Transmission Spectra

Abstract
Attempts to probe the atmospheres of rocky planets around M dwarfs present both promise and peril. While their favorable planet-to-star radius ratios enable searches for even thin secondary atmospheres, their high activity levels and high-energy outputs threaten atmosphere survival. Here, we present the 0.6–2.85 m transmission spectrum of the 1.1 , 340 K rocky planet TRAPPIST-1 c obtained over two JWST NIRISS/SOSS transit observations. Each of the two spectra displays 100–500 ppm signatures of stellar contamination. Despite being separated by 367 days, the retrieved spot and faculae properties are consistent between the two visits, resulting in nearly identical transmission spectra. Jointly retrieving for stellar contamination and a planetary atmosphere rules out with high confidence (3-) not only clear hydrogen-dominated atmospheres, but even thin, 1-bar high-mean molecular weight atmospheres rich in H2O, NH3, or CO (at the 2- level). We find that the only atmosphere scenarios which our spectrum cannot rule out are CH4- or CO2-rich atmospheres, which are both unlikely to be retained when considering the photodestruction of CH4 and the susceptibility of even a CO2-rich atmosphere to escape given the cumulative high-energy irradiation experienced by the planet. Our results further stress the importance of robustly accounting for stellar contamination when analyzing JWST observations of exo-Earths around M dwarfs, as well as the need for high-fidelity stellar models to search for the potential signals of thin secondary atmospheres.
1 Introduction
The detection and characterization of the atmospheres of rocky, Earth-sized planets is one of the key science goals of JWST. Not only is the presence of an atmosphere a top-level requirement for potential habitability (Grenfell, 2020), but constraining the composition of terrestrial planet atmospheres also has direct implications for our understanding of the formation and evolution of small planet atmospheres (Turbet et al., 2023; Krissansen-Totton & Fortney, 2022; Krissansen-Totton, 2023; Bolmont et al., 2017; Zahnle & Catling, 2017).
In recent years, the focus of such studies has been on small planets orbiting M dwarf stars, as the small size of these late-type host stars boosts the planet’s atmosphere signal, either in transmission or emission, compared to Sun-like stars (e.g., De Wit et al., 2016; Wakeford et al., 2019; Kreidberg et al., 2019; Lim et al., 2023; May et al., 2023; Greene et al., 2023; Zhang et al., 2024; Kirk et al., 2024). However, the potential promise of increased atmosphere observability also comes with a corresponding peril for atmosphere retention. Late-type stars have significantly longer pre-main sequence phases than Sun-like stars, during which time, orbiting planets are subjected to large amounts of high-energy radiation (Wordsworth & Pierrehumbert, 2013; Wheatley et al., 2017). Even at late times, during the M dwarf’s main sequence phase, they still maintain large outputs of high energy radiation, with ratios significantly higher than those of earlier-type stars (Pizzolato et al., 2003; Wright et al., 2011; Wheatley et al., 2017; Loyd et al., 2021), reaching up to 2000 that of the Sun for ultra-cool M dwarfs like TRAPPIST-1 (Wheatley et al., 2017). The increased amounts of high energy radiation received by M dwarf planets results in the potential for corresponding increases in the efficiency of atmosphere loss processes (e.g., Wordsworth & Pierrehumbert, 2013; Luger & Barnes, 2015; Bolmont et al., 2017; Turbet et al., 2023), resulting in uncertainty as to whether rocky planets around M dwarf stars can retain atmospheres to the present day (Zahnle & Catling, 2017).
Moreover, late-type stars also maintain higher levels of activity and photospheric inhomogeneities (i.e., the presence of spots and faculae) than Sun-like stars (Peacock et al., 2019; Loyd et al., 2021). Unocculted spots and faculae on the host star’s photosphere (i.e., inhomogeneities that lie outside of a planet’s transit chord), in particular, are a pernicious problem preventing the detection of M-Earth atmospheres via transit observations (Rackham et al., 2018, 2019; TRAPPIST-1 JWST Community Initiative et al., 2024). The presence of unocculted inhomogeneties, also known as the transit light source effect (TLSE), can impart spurious features in the transmission spectra of rocky planets, which can often be significantly larger than expected atmosphere features (Rackham et al., 2018, 2023; Lim et al., 2023; TRAPPIST-1 JWST Community Initiative et al., 2024). Despite these challenges, transit observations using the Hubble Space Telescope (HST) and now with JWST, have been able to widely rule out cloud-free, H/He-dominated atmospheres for rocky M dwarf planets (De Wit et al., 2016; Wakeford et al., 2019; Libby-Roberts et al., 2022; Lim et al., 2023; Lustig-Yaeger et al., 2023; Cadieux et al., 2024; Damiano et al., 2024; Kirk et al., 2024), although such extended atmospheres are already disfavoured based on mass and radius measurements alone (Agol et al., 2021). In many cases though, the TLSE, still prevents us from robustly probing secondary, or high-mean molecular weight atmospheres in transmission (e.g., Wakeford et al., 2019; Lim et al., 2023; Cadieux et al., 2024; May et al., 2023; Moran et al., 2023).
Observations of M-dwarf planets in eclipse geometry circumvent the TLSE problem but bring with them their own sets of challenges: the planet’s orbital eccentricity must be well known to calculate the phase of the eclipse, emission observations favour planets with high temperatures, potentially ruling out habitable-zone worlds, etc. Though, even without the impacts of the TLSE, emission studies of rocky M-dwarf planets, both in the JWST era and previously, have either been negative (Kreidberg et al., 2019; Greene et al., 2023; Ih et al., 2023; Zhang et al., 2024; Mansfield et al., 2024; Xue et al., 2024, i.e., favour atmosphereless interpretations;) or unable to provide definitive evidence for the presence of a high-mean molecular weight atmosphere (e.g., Zieba et al., 2023).
Here, we present transit observations of TRAPPIST-1 c with JWST NIRISS/SOSS. TRAPPIST-1 c is the second innermost planet (=0.01580 AU, =340 K) of the TRAPPIST-1 system, with a mass and radius of 1.3 and 1.1 that of the Earth, respectively (Gillon et al., 2016; Ducrot et al., 2020; Agol et al., 2021). It was first observed in transit by De Wit et al. (2016) with HST/WFC3 from 1.1 – 1.7 µm. Their observations rule out cloud-free, H/He-dominated atmosphere scenarios at 10-, but lack the precision to constrain the presence of a wide range (e.g., H2O-rich, Venus-like) of higher-mean molecular weight atmospheres. TRAPPIST-1 c was then observed in emission with JWST MIRI/F1500W photometry by Zieba et al. (2023) at 15 µm. Their dayside brightness temperature of 38031 K rules out thick (10 bar), CO2-rich atmospheres, but remains consistent with a thin (0.01 bar) atmosphere composed purely of CO2, or trace amounts of CO2 in thinner (0.1 – 10 bar), O2-dominated atmospheres at a 2 level.
This paper is organized as follows. We present the observations and data analysis in Section 2. We describe our analysis of the TLSE in Section 3, and joint modelling of the atmosphere and stellar contamination in Section 4. We then discuss our results in Section 5 and follow up with a short conclusion in Section 6.
2 Observations & Data Analysis
We observed two transits of TRAPPIST-1 c with the SOSS mode of JWST’s NIRISS instrument (Albert et al., 2023; Doyon et al., 2023) as part of program GO 2589 (PI: O. Lim). The first visit started at 18:37:51 UTC on October 28, 2022, and the second at 21:09:46 UTC on October 31, 2023. Each visit lasted a total of 4.6 hr, and consisted of 159 integrations with 18 groups per integration.
2.1 Data Reduction
We reduce the time series observations (TSO) with the exoTEDRF pipeline111https://github.com/radicamc/exoTEDRF, formerly known as supreme-SPOON. (Radica, 2024; Feinstein et al., 2023; Radica et al., 2023), closely following standard procedures (e.g., Radica et al., 2024; Cadieux et al., 2024). We perform the correction of 1/ noise at the group-level (that is, before ramp fitting) using the scale-achromatic-window method introduced in exoTEDRF v1.4.0. With this algorithm, we only use rows within 30 pixels of the trace to estimate the 1/ level, which can potentially help to reduce residual red noise in SOSS TSOs, particularly with high group numbers (e.g., Feinstein et al., 2023; Holmberg & Madhusudhan, 2023). We also perform a “piecewise” background subtraction (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024; Radica et al., 2024), whereby the standard STScI SOSS background model is scaled separately on either side of the background “step”, as we found that a single scaling for the whole detector did not adequately remove the background signal.
We extract the stellar spectrum using a simple box aperture with a width of 40 pixels, as the dilution effects due to the SOSS first and second order overlap are expected to be negligible for transit measurements (Darveau-Bernier et al., 2022; Radica et al., 2022). Finally, there are several undispersed contaminants due to background field stars (so-called “order 0” contaminants) present on the detector for both visits, several of which partially intersect the target spectral traces. These contaminants are located at wavelengths 0.670–0.675 µm, 1.642–1.658 µm, and 2.064–2.072 µm, for visit 1, and 0.975–1.001 µm, 1.251–1.279 µm, and 1.567–1.578 µm for visit 2. We simply mask these contaminants in the extracted spectra.
2.2 White Light Curve Fitting
We construct white light curves for each visit and each order by summing all flux on the detector in order 1 (0.85 – 2.8 µm), and only wavelengths 0.85 µm in order 2. We fit these white light curves using the flexible juliet package (Espinoza et al., 2019). For each visit, we jointly fit the first and second order white light curves, as in Radica et al. (2024). Concretely, we fit an astrophysical transit model, as computed by BATMAN (Kreidberg, 2015), with the orbital parameters (, the time of mid-transit, , the orbit impact parameter, and , the scaled orbital semi-major axis) shared between both orders, and the scaled planet radius, fit to each order individually. We, furthermore, assume a circular orbit and fix the orbital period to 2.421937 d (Agol et al., 2021). We fit, separately for each order, the two parameters of the quadratic limb darkening law, following the Kipping (2013) parameterization. We also fit a systematics model to the light curve for each order. For visit 2, the systematics model is composed of a linear trend with time, whereas for visit 1 we also include a quadratic term with time. TRAPPIST-1 is well-known to be a highly active star (Lim et al., 2023), and we, therefore, include a Gaussian process (GP) with a simple harmonic oscillator (SHO) kernel to model residual red noise in each light curve not captured by the aforementioned detrending. The SHO kernel, as implemented in celerite (Foreman-Mackey et al., 2017) has proven efficient at modelling the impacts of host star photospheric variations on light curves (i.e., stellar granulation, Kallinger et al., 2014; Pereira et al., 2019; Radica et al., 2024). For each visit, we fit the characteristic amplitude of the GP separately to each order but share the characteristic frequency between the two (e.g., Radica et al., 2024). Finally, we include an error inflation term for each order. Our final models therefore consist of 20 parameters for visit 1 and 18 for visit 2 — we use wide and uninformative priors in all cases.
The results of the white light curve fits are shown in Figure 1. During the pre-transit baseline of the second visit, there was a large flare event on the host star. We attempted several methods to correct this flare in the light curves e.g., detrending against the H flux, computed for each integration by measuring the area between the flux in the H line and a local quadratic continuum, shown in the second panels in Figure 1. However, this avenue was unsuccessful as there is a slight time delay between the appearance of the flare in the H time series and in the white light curve, likely due to lags in energy release as a function of wavelength (Howard et al., 2023). We eventually decided that, since the flare occurred in the pre-transit baseline and was sufficiently separated in time from the transit itself, the optimal course of action was to simply cut the affected integration from the fit. For visit 2, we, therefore, cut the first 75 integrations of the TSO (denoted with grey shading in Figure 1).
2.3 Spectroscopic Light Curve Fitting
To construct the spectrophotometric light curves, we sum the first order flux in bins of 80 pixels (corresponding to a constant resolving power of approximately =25). The combined effects of the spectrum of the late-M host star, and the low throughput of the SOSS order 2 mean that the second order trace is incredibly faint for observations of TRAPPIST-1 planets (e.g., Lim et al., 2023). We therefore do not bin order 2 further, and simply include the white light curve results in our transit spectrum. The orbital parameters from both visits are consistent between visits (Table 1), and we therefore fix them to the weighted average between the two when fitting the spectroscopically binned light curves. We include the same systematics model as mentioned above for each visit, except that we fix the characteristic timescale of the GP to the best-fitting value from the white light fit for each visit, and allow the amplitude to freely vary, since we can reasonably expect the impacts of stellar variability to be more prominent at bluer vs. redder wavelengths (e.g., Radica et al., 2024).
We find a slight offset of 122 ppm between the spectra of of the two visits, which we attribute to residual dilution from the pre-transit flare in visit 2. Indeed, this can also be noticed in the white light curve depths for each visit shown in Table 1, where the visit 2 order 1 transit depth is 100 ppm shallower than visit 1 — suggestive of dilution. We tested trimming a different number of pre-transit integrations and found that this slightly changed the offset (to the tune of 5 – 10 ppm), but did not change the shape of the final transit spectrum.
We also reduce and fit the TSOs from each visit using the independent NAMELESS pipeline (Feinstein et al., 2023; Coulombe et al., 2023). Further details of this analysis are presented in Appendix A, and a comparison of the spectra from both pipelines are shown in Figure 6. Since the spectra from each pipeline are consistent to within 1- for both visits, we only consider the exoTEDRF spectrum for the remainder of our analysis.
3 Modelling of the transit light source effect
We first model the transmission spectrum of TRAPPIST-1 c assuming that the TLSE can completely explain each visit’s transmission spectrum. This approach is motivated by the presence of small features resembling water absorption (e.g., around 1.4 µm; Figure 2) and a pronounced slope towards short wavelengths in the spectrum of the first visit. We use the open-source code stctm (Piaulet-Ghorayeb, 2024) to perform the TLS retrievals. We model the presence of two populations of heterogeneities: spots (covering a fraction of the photosphere, and cooler than the photosphere at a temperature ), and faculae (with representing their covering fraction, and their temperature).
For each population of heterogeneities, we fit for the temperature difference to the photosphere (), and place a conservative uniform prior from 0 to 50% on the covering fraction for each of the components, . The stellar photosphere temperature is fitted using a Gaussian prior, with the mean set by the literature value of 2566 K (Agol et al., 2021) and a standard deviation of 50 K. We allow faculae to be at most 1000 K hotter than the photosphere (with an otherwise uniform prior on ), while the minimum temperature of spots is set by the minimum temperature for which stellar models are available in our grid (2300 K). In the retrieval, we select at each iteration the closest-matching model (in effective temperature and surface gravity space) for the photosphere, spot, and faculae components from a finely-sampled pre-computed grid obtained with the MSG module (Townsend & Lopez, 2023) using the PHOENIX (Husser et al., 2013) stellar model grid. We also fit the used for the heterogeneities, allowing it to be lower than that of the photosphere following recent work (Lim et al., 2023; Fournier-Tondreau et al., 2024), but retrieve a value consistent with that of the photosphere (Agol et al., 2021).
We use the affine-invariant Markov Chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013) to explore the parameter space, with the number of walkers set as 20 times the number of fitted parameters, and run the chains for 5000 steps (60% are discarded as burn-in). In addition to the stellar heterogeneity properties, we fit the wavelength-independent transit depth of TRAPPIST-1 c, . We perform independent fits to the spectrum of each visit, as they are separated by about one year and should not, a priori, be affected by the same TLS contribution. We find that similar spot and facula properties can explain both visits (Figure 2).
4 Joint modelling of atmosphere and stellar contamination
We follow up our analysis of stellar contamination by using SCARLET (Benneke & Seager, 2012, 2013; Benneke, 2015; Benneke et al., 2019a, b; Pelletier et al., 2021; Piaulet et al., 2023) to more thoroughly explore the potential atmosphere scenarios consistent with our observation, while still accounting for the contributions of stellar contamination. The forward modelling component of SCARLET iteratively solves the radiative transfer equation and the 1D vertical structure of the atmosphere under the assumption of hydrostatic equilibrium, given the vertical temperature-pressure profile and the molecular abundances in each atmospheric layer. In the retrieval, the planetary radius is optimized for each set of sampled parameters to obtain the best match to the observed spectrum.
We perform three categories of retrievals corresponding to different assumptions about the planetary atmosphere: (1) H2/He-dominated scenarios; (2) single-component high-mean molecular weight atmospheres; (3) two-component high-mean molecular weight atmospheres with one infrared absorber present in the atmosphere and either N2 or O2 acting as the background gas.
We parameterize H2/He-dominated atmospheres in the retrieval with a H2/He mixture as the background gas (assuming He/H2= 0.157), log-uniform priors on the abundances of N2, CH4, H2O, CO, CO2, NH3, and SO2, and a log-uniform prior on the pressure at which the atmosphere becomes opaque (either due to the presence of a cloud deck or a solid surface).
For single-component high-mean molecular weight atmospheres, we only explore absorbers with prominent features over the NIRISS/SOSS wavelength range; namely, atmospheres made of 100% CH4, H2O, CO, or NH3. We also fit for the “effective surface pressure” at which the atmosphere becomes opaque.
Finally, we also explore potential N2-dominated atmospheres (similar to some terrestrial objects in the solar system), or O2-dominated atmospheres (motivated by the possible O2-CO2 atmosphere consistent with the MIRI eclipse observations of TRAPPIST-1 c; (Zieba et al., 2023)) where only one of CH4, H2O, CO, or NH3 acts as an absorber, beyond the broad N2-N2 or O2-O2 collision-induced absorption features introduced by the background gas. In this retrieval setup, we fit for the partial pressure of each component.
We assume that the vertical temperature structure is isothermal, motivated by the likely-thin region of the upper atmosphere probed by our observations in transmission, and only fit for this photosphere temperature. We use the nested sampling method (Skilling, 2004, 2006) to sample the full parameter space, with the multi-ellipsoid method implemented using the nestle module 222https://github.com/kbarbary/nestle within the SCARLET retrieval framework. The final posterior distributions are obtained from retrievals performed using at least 1000 live points. We compute forward models at R=15,625 before convolving them to the resolution of the data assuming a uniform throughput within each bin for the likelihood calculation.
We also calculate forward models for 1 and 100 solar metallicity, H2/He-dominated atmospheres with varying cloud-top pressures (Figure 3; left panel), and models for several high mean molecular weight compositions (Figure 4; top panel), to put our retrieval results in context. For each model, the temperature-pressure profile is assumed to be isothermal at the equilibrium temperature of TRAPPIST-1 c.
For each set of retrievals, we account in the modelling for the fact that the transmission spectra we observe are likely heavily affected by stellar contamination (see Figure 2), thereby marginalizing over the interplay that unocculted stellar heterogeneities can have with planetary absorption features. In addition to the planetary atmosphere properties, we thus account for stellar heterogeneities assuming a single population of spots, fitting for the spot covering fraction and their temperature contrast with the photosphere, as well as the photosphere temperature itself. We also assume that the stellar heterogeneity components of both visits are identical, motivated by the results from the stctm fit (Figure 2), and do not fit for a different for the heterogeneity component. We note, though, that we performed one set of atmosphere+stellar contamination retrievals where we fit for visit-specific stellar heterogeneity properties, and found that this does not affect our inferences, further motivating our choice of shared spot parameters for both visits. We place uniform priors on the spot fraction and temperature, but use the same Gaussian prior as in the TLS-only fit for the photosphere temperature.
We note that we find a vertical offset of 100 ppm between the spectra of both visits (as can be inferred from the white light transit radii presented in Table 1). We attribute it to the presence of the flare in the pre-transit baseline of the second visit that may result in a dilution effect when fitting the transit light curves and therefore fit for a parameter describing the visit 1-visit 2 offset.
5 Discussion
5.1 Impacts of Stellar Contamination
Our two-transit study of TRAPPIST-1 c aptly summarizes the promise and peril of transmission observations of rocky planets around M dwarf host stars. There are now a plethora of studies that have shown that stellar contamination is a true bottleneck in the attempt to detect atmospheres around rocky, M dwarf planets via transmission spectroscopy (e.g., May et al., 2023; Moran et al., 2023; Lim et al., 2023; Cadieux et al., 2024). In most cases, different manifestations of the TLSE (i.e., different temperature distributions, covering fractions, etc. of stellar spots and faculae) give rise to obviously different transmission spectra between visits, allowing for a clear inference of the impacts of the TLSE, even without undertaking in-depth modelling. Consistency in the location and shapes of spectral features between visits has been argued to lend credence to an interpretation favouring a planetary atmosphere origin (e.g. May et al., 2023).
However, despite the high level of consistency between the transmission spectra (1.1- on average), our retrieval analysis still finds that both visits can be entirely explained by stellar contamination. As shown in Figure 2, we find very similar stellar heterogeneity parameters (spot/faculae temperatures and covering fractions) for both visits. The spot and faculae temperatures are consistent with those found by Lim et al. (2023) in their analysis of NIRISS/SOSS transmission spectra of TRAPPIST-1 b. However, we find both visits to be spot-dominated, with minimal contributions from faculae, whereas Lim et al. (2023) have one spot-dominated, and one facula-dominated visit.
The consistency between visits would perhaps be unsurprising if they were separated by a very small period of time, which would therefore probe similar spot and facula distributions on the host star. However, our two visits are separated by 367 d, significantly longer than the 3.3 d stellar rotation period (Luger et al., 2017). When phase-folded to the stellar rotation period, this corresponds to a phase difference of 0.21 — similar to the observations of Lim et al. (2023), who, as mentioned above, found no such consistency between their two visits. This highlights the fact that TLSE models are generally non-predictive of future stellar contamination, since the distributions of spots and faculae in the host star’s photosphere is entirely stochastic (Rackham et al., 2018, 2023; TRAPPIST-1 JWST Community Initiative et al., 2024).
5.2 Cloud-Free and Thick Hydrogen-Rich Atmospheres Are Ruled Out
We find that, similarly to the closer-in planet TRAPPIST-1 b (Lim et al., 2023), we can confidently rule out thick or cloud-free hydrogen-dominated atmospheres, with a 3- upper limit of only 0.1 bar on the effective surface pressure of a H2/He-dominated atmosphere atop TRAPPIST-1 c (Figure 3). The thin 0.1 bar hydrogen-rich atmosphere scenarios allowed by the transmission spectrum are, moreover, unlikely from an atmospheric evolution standpoint — TRAPPIST-1 c was exposed to large cumulative amounts of high-energy XUV irradiation from its host star, sufficient to strip away any primordial atmosphere (Hori & Ogihara, 2020; Turbet et al., 2020) even when considering hydrogen outgassing (Hu et al., 2023). The transmission spectrum still allows within the 3- confidence contours for thick, high-metallicity (100 solar) atmospheres, which may be more resilient against escape due to their higher mean molecular weights.
5.3 Limits on High Mean Molecular Weight Atmospheres
We, therefore, turn to evaluate the range of plausible high-metallicity atmosphere scenarios for TRAPPIST-1 c, and focus on the detectability of H2O-, NH3-, CO- and CH4-rich atmospheres, as these infrared absorbers have sufficient opacity within the NIRISS/SOSS wavelength range to obtain meaningful constraints. We can rule out thick (1 bar) 100% H2O and NH3 atmospheres at better than 3, and even a 1 bar 100% CO atmosphere is excluded at the 2- level, although thinner 0.1 bar atmospheres remain plausible explanations for the flat spectrum (Figure 4).
With the exception of CH4, we find tight upper limits on the abundance of all the other molecules tested (H2O, NH3, CO), even in lower-mean molecular weight atmospheres dominated by N2 or O2 as the background gas (Figure 5). We compare those limits, where available, with measured compositions of solar system planets. In particular, we highlight in Figure 5 the amount of H2O in Earth’s N2-dominated atmosphere (using 0.4% as the mean mole fraction over the entire 1-bar atmosphere, although it can reach about one percent at sea level; Kasai et al., 2011; Kelsey et al., 2022), or the amount of CH4 in Titan’s atmosphere, assuming either a cloud-free Titan scenario (where one can see down to the surface) with 5% CH4, or a hazy scenario accounting for the fact that hydrocarbon haze opacity raises the effective surface pressure to 10 mbar, and with a factor of three less methane in the upper atmosphere (see e.g. Charnay et al., 2014. We find that similar amounts of water to what is observed in Earth’s atmosphere are ruled out for TRAPPIST-1 c, while the transmission spectrum alone is not sensitive enough to the CH4 abundance to offer meaningful constraints on Titan-like atmospheres. Although our observations cannot rule out the presence of CH4 either in thick 1–100 bar pure-CH4 atmospheres, or as a trace species in a N2- or O2-dominated atmosphere, CH4 is unlikely on TRAPPIST-1 c from a modelling standpoint because of its vulnerability to destruction via photochemistry (Turbet et al., 2018).
Since O2 is a heavier molecule than N2, the O2 background scenarios generally correspond to more compact, higher mean molecular weight atmospheres. Therefore, although the results are qualitatively similar to the N2-background constraints, our transmission spectrum allows for larger amounts of the tested molecular absorbers in an O2-dominated atmosphere. As such, we unfortunately are not able to rule on any of the possible scenarios presented by Zieba et al. (2023) based on their MIRI eclipse observations of this planet. The photometric eclipse depth of 42194 ppm at 15 µm was broadly consistent (better than 2) with a 0.01 bar pure CO2 atmosphere, or trace amounts of CO2 in a O2-dominated atmospheres with a surface pressure less than 10 bar. We find that such an O2-CO2 atmosphere would give rise in transmission to a 40 ppm CO2 absorption feature at 2.7 µm (at the red edge of the SOSS detector), and a larger 90 ppm absorption band at 4.3 µm. As demonstrated by Taylor et al. (2023), CO2 cannot be reliably probed at 2.7 µm with SOSS alone, even for a hot-Jupiter, whereas the 4.3 µm feature could potentially be detectable with five transits with NIRSpec/G395M or PRISM.
5.4 Implications for the Presence of an Atmosphere on TRAPPIST-1c
Ultimately, our results from the analysis of the observed transmission spectrum of TRAPPIST-1 c are in line with modeling predictions that it would not have been able to retain any substantial atmosphere over the 5–9 Gyr lifetime of its host star (Burgasser & Mamajek, 2017), regardless of its initial volatile content (Krissansen-Totton, 2023).
The thin, 1 bar atmospheres that remain allowed by our observations, even if they correspond to high mean molecular weights, are unlikely to be retained on TRAPPIST-1 c from a theoretical perspective. Even if Jeans escape (generally less efficient than blow-off) is expected to dominate for high-mean molecular weight atmospheres atop TRAPPIST-1 c, the escape rates even for resilient N2-CO2 atmospheres range from one Earth atmosphere per Myr to one Earth atmosphere mass per Gyr (Van Looveren et al., 2024) at the present-day irradiation of this planet. The implications are even stricter when considering the fact that any atmosphere on TRAPPIST-1 c also would have had to withstand the early XUV-active stages of the late M dwarf’s evolution (Bolmont et al., 2017), and its position inside of the runaway greenhouse limit (Turbet et al., 2020) indicates that water could not have been efficiently shielded from escape via condensation in a liquid water ocean, even if this limit is pushed to closer host star-planet distances around M dwarfs in 3D models (Turbet et al., 2023).
6 Conclusions
Here, we presented the first JWST transmission spectrum of the second innermost planet in the TRAPPIST-1 system; TRAPPIST-1 c. Our 0.6 – 2.85 µm NIRISS/SOSS spectrum reveals strong contamination by the TLSE in both visits. However, unlike in previous multi-visit analyses of rocky planets around M dwarfs (e.g., May et al., 2023; Moran et al., 2023; Cadieux et al., 2024), including TRAPPIST-1 b (Lim et al., 2023), we find consistent stellar heterogeneity properties between our two visits, resulting in nearly identical manifestations of the TLSE. Our results support the need for joint modelling of the TLSE and atmosphere properties to achieve robust constraints on rocky, M dwarf planets in the era of JWST. Importantly, further stellar model fidelity is critically required to accurately model out the impact of the TLSE and reach the sensitivity to detect or rule out thin atmospheres on small planets via their transit spectra.
When accounting for stellar contamination in our atmosphere analysis, we find that thick, clear hydrogen-dominated compositions are ruled out at better than 3- significance, and even high-mean molecular weight compositions rich in H2O, NH3 or CO are excluded for an atmosphere thickness of 1 bar or more at better than 2. The NIRISS/SOSS transmission spectrum of TRAPPIST-1 c alone cannot rule out all high mean molecular weight compositions, especially for CH4- and CO2-rich cases, to which our spectrum is not highly sensitive. However, in light of modelling predictions that CH4-rich compositions are unlikely as CH4 is sensitive to photochemical destruction (Turbet et al., 2018), and considering that Jeans escape would quickly strip TRAPPIST-1 c of even CO2-rich atmospheres (Van Looveren et al., 2024), we conclude that this planet is unlikely to bear any atmosphere whatsoever.
Regardless, as shown by Krissansen-Totton (2023), even if TRAPPIST-1 c is completely atmosphere-less, this does not imply a similar fate for the outer, habitable-zone planets. Our work, though, does highlight the fact that we cannot simply rely on consistency of transmission spectra between different visits to infer the presence of a planetary atmosphere and rule out stellar contamination.
Appendix A NAMELESS Reduction and ExoTEP Light Curve Fitting
We reduce our two JWST NIRISS/SOSS visits of TRAPPIST-1c, starting from the uncalibrated raw data up to the extraction of the stellar spectra, using the NAMELESS pipeline (Coulombe et al., 2023; Feinstein et al., 2023). We first go through the Stages 1 and 2 steps of the STScI jwst pipeline (v1.12.5) (Bushouse et al., 2022), which include super-bias subtraction, reference pixel correction, non-linearity correction, ramp-fitting, and flat-fielding. We then proceed with a sequence of custom steps to correct for remaining sources of noise such as bad pixels, non-uniform background, cosmic-rays, and 1/ following the methods described in Benneke et al. (2024). We flag bad pixels by looking for pixels whose second derivative are significant outliers in the spatial domain. The counts of all flagged pixels are then corrected using the bicubic interpolation method of the scipy.interpolate.griddata function. We correct for the non-uniform background by individually scaling the two regions of the STScI model background333https://jwst-docs.stsci.edu/ that are separated by the sudden jump in background flux situated around spectral pixel x700. Any remaining cosmic rays are corrected by clipping any count that is more than four standard deviations away from the running median for all pixels. As for the 1/ noise correction, we follow the same method described in Coulombe et al. (2023) and Benneke et al. (2024) where we scale individually each column of the first and second spectral orders considering only pixels that are within a 30-pixel distance from the center of the traces to compute the 1/ noise. We then extract the stellar spectra from both spectral orders using a simple box aperture with a width of 36 pixels.
We perform the broadband and spectroscopic fits of the light curves produced by the NAMELESS reduction with the ExoTEP framework (Benneke et al., 2019a, b). For our fit of the white-light curve, which is obtained by summing all wavelengths of the first spectral order (0.85–2.85 m), we keep free the mid-transit time (, considering uniform priors spanning the time of the observations for each visit), planet-to-star radius ratio (, [0.01,0.2]), semi-major axis (, [15,45]), and the impact parameter (, [0,1]). We consider the quadratic limb-darkening law and fit directly for and assuming large uniform priors ([-3,3]) following Coulombe et al. (2024). For the first visit, we consider a quadratic trend for the systematics model. We also include in the fit a Matérn 3/2 Gaussian Process using the celerite python (Foreman-Mackey et al., 2017) to remove higher-frequency trends that are not corrected by our systematics model. The timescale and amplitude of the GP model are kept free assuming large uninformative uniform priors in log-space. As for the second visit, we cut the first 2.1 hours of the TSO and consider solely a linear slope for the systematics model. The transit light curves are then modelled in ExoTEP using the batman (Kreidberg, 2015) python package, and we explore the parameter space using and the Markov chain Monte Carlo (MCMC) sampler emcee (Foreman-Mackey et al., 2013). We run the MCMC for 10,000 steps and discard the first 6,000 steps as burn-in. The same procedure is repeated for the spectroscopic light curves (assuming the same systematics model and GP for each visit), where we have fixed the orbital parameters of TRAPPIST-1 c to the best-fit values from the white-light curve of the first visit ( = 28.55, = 0.112). The light curves are fit using 100 and 30 equal-pixel bins for the first and second spectral orders, respectively. The resulting spectra, binned to a fixed resolving power of , are shown in Figure 6.
Appendix B Additional Tables
Table 1 contains the best-fitting white light curve parameters for both NIRISS/SOSS visits. The orbital parameters (, ) are consistent with each other, and with previous literature (e.g., Agol et al., 2021).
| Parameter | Value | ||
| Visit 1 | Visit 2 | Weighted Average | |
| T0 [BJD] | 2459881.401521 | 2460249.515096 | - |
| 0.133 | 0.136 | ||
| 28.31 | 28.12 | ||
Notes: For each visit, and the limb darkening parameters ( and ) were fit separately for each order. All other parameters were jointly fit to both.
Table 2 has the exoTEDRF transmission spectrum used in this work.
| Wavelength | Wavelength Error | Transit Depth (visit 1) | Transit Depth (visit 2) |
| (µm) | (µm) | (ppm) | (ppm) |
| 0.72 | 0.12 | 6050.99 473.23 | 6982.25 418.63 |
| 0.90 | 0.04 | 6585.89 111.83 | 6757.96 135.93 |
| 0.98 | 0.04 | 6772.63 91.97 | 6808.19 128.63 |
| 1.05 | 0.04 | 6881.03 71.42 | 6960.16 89.30 |
| 1.12 | 0.04 | 6820.05 69.73 | 6849.88 64.50 |
| 1.20 | 0.04 | 6699.58 63.55 | 6907.84 80.39 |
| 1.28 | 0.04 | 6694.65 64.58 | 6697.34 90.68 |
| 1.35 | 0.04 | 6971.35 60.31 | 7023.65 58.31 |
| 1.43 | 0.04 | 6886.42 76.10 | 7001.35 76.95 |
| 1.51 | 0.04 | 6847.72 84.78 | 6937.97 88.64 |
| 1.59 | 0.04 | 6789.90 76.61 | 6902.63 90.38 |
| 1.66 | 0.04 | 6895.75 74.07 | 7088.25 92.70 |
| 1.74 | 0.04 | 6914.99 90.90 | 6899.64 100.01 |
| 1.82 | 0.04 | 7006.82 97.16 | 7136.91 144.17 |
| 1.90 | 0.04 | 6804.84 126.15 | 6967.40 122.15 |
| 1.98 | 0.04 | 7032.96 118.65 | 7127.79 126.21 |
| 2.06 | 0.04 | 6796.19 132.58 | 6970.70 136.08 |
| 2.14 | 0.04 | 6842.86 136.57 | 6972.46 134.27 |
| 2.21 | 0.04 | 7010.63 135.26 | 6977.26 147.22 |
| 2.29 | 0.04 | 6919.63 128.65 | 6890.04 155.09 |
| 2.37 | 0.04 | 6707.69 153.51 | 6972.92 152.57 |
| 2.45 | 0.04 | 6643.73 180.88 | 6821.97 202.91 |
| 2.53 | 0.04 | 6770.44 232.33 | 7296.84 270.63 |
| 2.61 | 0.04 | 6837.59 233.79 | 6966.78 293.63 |
| 2.69 | 0.04 | 6183.53 235.11 | 6961.30 273.61 |
| 2.77 | 0.04 | 6589.56 270.29 | 6630.95 311.46 |

### v3

@@
 Despite being separated by 367 days, the retrieved spot and faculae properties are consistent between the two visits, resulting in nearly identical transmission spectra.
-Jointly retrieving for stellar contamination and a planetary atmosphere rules out with high confidence (3-) not only clear hydrogen-dominated atmospheres, but even thin, 1-bar high-mean molecular weight atmospheres rich in H2O, NH3, or CO (at the 2- level).
-We find that the only atmosphere scenarios which our spectrum cannot rule out are CH4- or CO2-rich atmospheres, which are both unlikely to be retained when considering the photodestruction of CH4 and the susceptibility of even a CO2-rich atmosphere to escape given the cumulative high-energy irradiation experienced by the planet.
+Jointly retrieving for stellar contamination and a planetary atmosphere reveals that our spectrum can rule out hydrogen-dominated, 300 solar metallicity atmospheres with effective surface pressures down to 10 mbar at the 3- level.
+For high-mean molecular weight atmospheres, where O2 or N2 is the background gas, our spectrum disfavors partial pressures of more than 10 mbar for H2O, CO, NH3 and CH4 at the 2- level.
+Similarly, under the assumption of a 100% H2O, NH3, CO, or CH4 atmosphere, our spectrum disfavors thick, 1 bar atmospheres at the 2- level.
+These non-detections of spectral features are in line with predictions that even heavier, CO2-rich, atmospheres would be efficiently lost on TRAPPIST-1 c given the cumulative high-energy irradiation experienced by the planet.
 Our results further stress the importance of robustly accounting for stellar contamination when analyzing JWST observations of exo-Earths around M dwarfs, as well as the need for high-fidelity stellar models to search for the potential signals of thin secondary atmospheres.
@@
 Not only is the presence of an atmosphere a top-level requirement for potential habitability (Grenfell, 2020), but constraining the composition of terrestrial planet atmospheres also has direct implications for our understanding of the formation and evolution of small planet atmospheres (Turbet et al., 2023; Krissansen-Totton & Fortney, 2022; Krissansen-Totton, 2023; Bolmont et al., 2017; Zahnle & Catling, 2017).
-In recent years, the focus of such studies has been on small planets orbiting M dwarf stars, as the small size of these late-type host stars boosts the planet’s atmosphere signal, either in transmission or emission, compared to Sun-like stars (e.g., De Wit et al., 2016; Wakeford et al., 2019; Kreidberg et al., 2019; Lim et al., 2023; May et al., 2023; Greene et al., 2023; Zhang et al., 2024; Kirk et al., 2024).
+In recent years, the focus of such studies has been on small planets orbiting M dwarf stars, as the small size of late-type stars boosts the planet’s atmosphere signal, either in transmission or emission, compared to Sun-like stars (e.g., De Wit et al., 2016; Wakeford et al., 2019; Kreidberg et al., 2019; Lim et al., 2023; May et al., 2023; Greene et al., 2023; Zhang et al., 2024; Kirk et al., 2024).
 However, the potential promise of increased atmosphere observability also comes with a corresponding peril for atmosphere retention.
-Late-type stars have significantly longer pre-main sequence phases than Sun-like stars, during which time, orbiting planets are subjected to large amounts of high-energy radiation (Wordsworth & Pierrehumbert, 2013; Wheatley et al., 2017).
-Even at late times, during the M dwarf’s main sequence phase, they still maintain large outputs of high energy radiation, with ratios significantly higher than those of earlier-type stars (Pizzolato et al., 2003; Wright et al., 2011; Wheatley et al., 2017; Loyd et al., 2021), reaching up to 2000 that of the Sun for ultra-cool M dwarfs like TRAPPIST-1 (Wheatley et al., 2017).
+Late-type stars have significantly longer pre-main sequence phases than Sun-like stars, during which time orbiting planets are subjected to large amounts of high-energy radiation (Wordsworth & Pierrehumbert, 2013; Wheatley et al., 2017).
+Even at late times, during the M dwarf’s main sequence phase, they still maintain large outputs of high energy radiation, with ratios significantly higher than those of earlier-type stars (Pizzolato et al., 2003; Wright et al., 2011; Wheatley et al., 2017; Loyd et al., 2021), reaching up to 2000 that of the Sun for ultra-cool M dwarfs like TRAPPIST-1 (Wheatley et al., 2017; Bourrier et al., 2017).
 The increased amounts of high energy radiation received by M dwarf planets results in the potential for corresponding increases in the efficiency of atmosphere loss processes (e.g., Wordsworth & Pierrehumbert, 2013; Luger & Barnes, 2015; Bolmont et al., 2017; Turbet et al., 2023), resulting in uncertainty as to whether rocky planets around M dwarf stars can retain atmospheres to the present day (Zahnle & Catling, 2017).
 Moreover, late-type stars also maintain higher levels of activity and photospheric inhomogeneities (i.e., the presence of spots and faculae) than Sun-like stars (Peacock et al., 2019; Loyd et al., 2021).
-Unocculted spots and faculae on the host star’s photosphere (i.e., inhomogeneities that lie outside of a planet’s transit chord), in particular, are a pernicious problem preventing the detection of M-Earth atmospheres via transit observations (Rackham et al., 2018, 2019; TRAPPIST-1 JWST Community Initiative et al., 2024).
-The presence of unocculted inhomogeneties, also known as the transit light source effect (TLSE), can impart spurious features in the transmission spectra of rocky planets, which can often be significantly larger than expected atmosphere features (Rackham et al., 2018, 2023; Lim et al., 2023; TRAPPIST-1 JWST Community Initiative et al., 2024).
-Despite these challenges, transit observations using the Hubble Space Telescope (HST) and now with JWST, have been able to widely rule out cloud-free, H/He-dominated atmospheres for rocky M dwarf planets (De Wit et al., 2016; Wakeford et al., 2019; Libby-Roberts et al., 2022; Lim et al., 2023; Lustig-Yaeger et al., 2023; Cadieux et al., 2024; Damiano et al., 2024; Kirk et al., 2024), although such extended atmospheres are already disfavoured based on mass and radius measurements alone (Agol et al., 2021).
+Unocculted spots and faculae on the host star’s photosphere (i.e., inhomogeneities that lie outside of a planet’s transit chord), in particular, are a pernicious problem preventing the detection of atmospheres around Earth-like planets orbiting M dwarfs via transit observations (Rackham et al., 2018, 2019; Wakeford et al., 2019; Garcia et al., 2022; TRAPPIST-1 JWST Community Initiative et al., 2024; Rackham & de Wit, 2024).
+The presence of unocculted inhomogeneities can impart spurious features on the transmission spectra of rocky planets, which can often be significantly larger than expected atmosphere features (Rackham et al., 2018, 2023; Lim et al., 2023; TRAPPIST-1 JWST Community Initiative et al., 2024; Rackham & de Wit, 2024).
+This phenomenon is often referred to as the transit light source effect (TLSE).
+Despite these challenges, transit observations using the Hubble Space Telescope (HST) and now with JWST, have been able to widely rule out cloud-free, H/He-dominated atmospheres for rocky M dwarf planets (De Wit et al., 2016; Wakeford et al., 2019; Garcia et al., 2022; Libby-Roberts et al., 2022; Lim et al., 2023; Lustig-Yaeger et al., 2023; Cadieux et al., 2024; Damiano et al., 2024; Kirk et al., 2024), although such extended atmospheres are often already disfavoured based on mass and radius measurements alone (e.g., Agol et al., 2021).
 In many cases though, the TLSE, still prevents us from robustly probing secondary, or high-mean molecular weight atmospheres in transmission (e.g., Wakeford et al., 2019; Lim et al., 2023; Cadieux et al., 2024; May et al., 2023; Moran et al., 2023).
-Observations of M-dwarf planets in eclipse geometry circumvent the TLSE problem but bring with them their own sets of challenges: the planet’s orbital eccentricity must be well known to calculate the phase of the eclipse, emission observations favour planets with high temperatures, potentially ruling out habitable-zone worlds, etc.
-Though, even without the impacts of the TLSE, emission studies of rocky M-dwarf planets, both in the JWST era and previously, have either been negative (Kreidberg et al., 2019; Greene et al., 2023; Ih et al., 2023; Zhang et al., 2024; Mansfield et al., 2024; Xue et al., 2024, i.e., favour atmosphereless interpretations;) or unable to provide definitive evidence for the presence of a high-mean molecular weight atmosphere (e.g., Zieba et al., 2023).
+Observations of M dwarf planets in eclipse geometry circumvent the TLSE problem but bring with them their own sets of challenges: the planet’s orbital eccentricity must be well known to calculate the phase of the eclipse, emission observations favour planets with high temperatures potentially ruling out habitable-zone worlds, etc.
+Though, even without the impacts of the TLSE, emission studies of rocky M-dwarf planets, both in the JWST era and previously, have either been negative (Kreidberg et al., 2019; Greene et al., 2023; Ih et al., 2023; Zhang et al., 2024; Mansfield et al., 2024; Xue et al., 2024, i.e., favour atmosphereless interpretations;) or unable to provide definitive evidence for the presence of a high-mean molecular weight (MMW) atmosphere (e.g., Zieba et al., 2023; Lincowski et al., 2023), often due to degeneracies between atmospheric absorption and surface reflectivity (e.g., Ih et al., 2023; Lincowski et al., 2023).
 Here, we present transit observations of TRAPPIST-1 c with JWST NIRISS/SOSS.
@@
 (2016) with HST/WFC3 from 1.1 – 1.7 µm.
-Their observations rule out cloud-free, H/He-dominated atmosphere scenarios at 10-, but lack the precision to constrain the presence of a wide range (e.g., H2O-rich, Venus-like) of higher-mean molecular weight atmospheres.
+Their observations rule out cloud-free, H/He-dominated atmosphere scenarios at 10-, but lack the precision to constrain the presence of a wide range (e.g., H2O-rich, Venus-like) of higher-MMW atmospheres.
 TRAPPIST-1 c was then observed in emission with JWST MIRI/F1500W photometry by Zieba et al.
@@
 Their dayside brightness temperature of 38031 K rules out thick (10 bar), CO2-rich atmospheres, but remains consistent with a thin (0.01 bar) atmosphere composed purely of CO2, or trace amounts of CO2 in thinner (0.1 – 10 bar), O2-dominated atmospheres at a 2 level.
+This was then followed up by Lincowski et al.
+(2023) who compared the Zieba et al.
+(2023) eclipse depth to a broader range of potential atmosphere scenarios using a self-consistent modelling framework.
+Their results generally agree with those of Zieba et al.
+(2023), in that trace amounts of CO2 in a thin O2-dominated atmosphere provide the best matches to the data.
+However, they also also find that 0.1 bar pure-O2, and even 3 bar steam atmospheres are consistent with the data within 2.
+Venus-like atmospheres, though, are disfavoured at 3.
 This paper is organized as follows.
 We present the observations and data analysis in Section 2.
-We describe our analysis of the TLSE in Section 3, and joint modelling of the atmosphere and stellar contamination in Section 4.
+We describe our analysis of the stellar photosphere in Section 3, and joint modelling of the atmosphere and stellar contamination in Section 4.
 We then discuss our results in Section 5 and follow up with a short conclusion in Section 6.
@@
 2.1 Data Reduction
-We reduce the time series observations (TSO) with the exoTEDRF pipeline111https://github.com/radicamc/exoTEDRF, formerly known as supreme-SPOON.
-(Radica, 2024; Feinstein et al., 2023; Radica et al., 2023), closely following standard procedures (e.g., Radica et al., 2024; Cadieux et al., 2024).
-We perform the correction of 1/ noise at the group-level (that is, before ramp fitting) using the scale-achromatic-window method introduced in exoTEDRF v1.4.0.
+We reduce the time series observations (TSO) with the exoTEDRF pipeline111https://github.com/radicamc/exoTEDRF (Radica, 2024; Feinstein et al., 2023; Radica et al., 2023), closely following standard procedures as described in, e.g., Radica et al.
+(2024); Cadieux et al.
+(2024); Piaulet-Ghorayeb et al.
+(2024).
+Concretely, we perform the standard exoTEDRF stage 1 calibrations, including the correction of 1/ noise at the group-level (that is, before ramp fitting) using the scale-achromatic-window method introduced in exoTEDRF v1.4.0.
 With this algorithm, we only use rows within 30 pixels of the trace to estimate the 1/ level, which can potentially help to reduce residual red noise in SOSS TSOs, particularly with high group numbers (e.g., Feinstein et al., 2023; Holmberg & Madhusudhan, 2023).
-We also perform a “piecewise” background subtraction (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024; Radica et al., 2024), whereby the standard STScI SOSS background model is scaled separately on either side of the background “step”, as we found that a single scaling for the whole detector did not adequately remove the background signal.
+We then proceed with the standard Stage 2 calibrations (e.g., flat field correction, bad pixel interpolation, etc.).
+Notably, we perform a “piecewise” background subtraction (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024; Radica et al., 2024), whereby the standard STScI SOSS background model is scaled separately on either side of the background “step”, as we found that a single scaling for the whole detector did not adequately remove the background signal.
+As a final step in the calibrations, we performed a principle component analysis (PCA) on the 2D detector images (e.g., Coulombe et al., 2023; Radica et al., 2024) to identify any detector-level trends (i.e., mirror tilt events, trace position drifts), which could impact the light curves.
 We extract the stellar spectrum using a simple box aperture with a width of 40 pixels, as the dilution effects due to the SOSS first and second order overlap are expected to be negligible for transit measurements (Darveau-Bernier et al., 2022; Radica et al., 2022).
@@
 We construct white light curves for each visit and each order by summing all flux on the detector in order 1 (0.85 – 2.8 µm), and only wavelengths 0.85 µm in order 2.
-We fit these white light curves using the flexible juliet package (Espinoza et al., 2019).
+We fit these white light curves using the flexible exoUPRF library (Radica, 2024).
 For each visit, we jointly fit the first and second order white light curves, as in Radica et al.
 (2024).
-Concretely, we fit an astrophysical transit model, as computed by BATMAN (Kreidberg, 2015), with the orbital parameters (, the time of mid-transit, , the orbit impact parameter, and , the scaled orbital semi-major axis) shared between both orders, and the scaled planet radius, fit to each order individually.
+Concretely, we fit an astrophysical transit model, as computed by BATMAN (Kreidberg, 2015), with the orbital parameters (, the time of mid-transit, , the orbital inclination, and , the scaled orbital semi-major axis) shared between both orders, and the scaled planet radius, fit to each order individually.
 We, furthermore, assume a circular orbit and fix the orbital period to 2.421937 d (Agol et al., 2021).
-We fit, separately for each order, the two parameters of the quadratic limb darkening law, following the Kipping (2013) parameterization.
+We fit, separately for each order, the two parameters of the quadratic limb darkening law, , in order to avoid potential biases that can be introduced by employing the Kipping (2013), parameterization (e.g., Coulombe et al., 2024).
+We also tested several other limb-darkening treatments but found that the choice of parameterization did not impact the final spectra.
 We also fit a systematics model to the light curve for each order.
-For visit 2, the systematics model is composed of a linear trend with time, whereas for visit 1 we also include a quadratic term with time.
-TRAPPIST-1 is well-known to be a highly active star (Lim et al., 2023), and we, therefore, include a Gaussian process (GP) with a simple harmonic oscillator (SHO) kernel to model residual red noise in each light curve not captured by the aforementioned detrending.
-The SHO kernel, as implemented in celerite (Foreman-Mackey et al., 2017) has proven efficient at modelling the impacts of host star photospheric variations on light curves (i.e., stellar granulation, Kallinger et al., 2014; Pereira et al., 2019; Radica et al., 2024).
-For each visit, we fit the characteristic amplitude of the GP separately to each order but share the characteristic frequency between the two (e.g., Radica et al., 2024).
-Finally, we include an error inflation term for each order.
-Our final models therefore consist of 20 parameters for visit 1 and 18 for visit 2 — we use wide and uninformative priors in all cases.
-The results of the white light curve fits are shown in Figure 1.
+For each visit, the first component of the systematics model is a linear trend with time.
+TRAPPIST-1 is well-known to be a highly active star (Lim et al., 2023), and we, therefore, include a Gaussian process (GP) with a Matérn 3/2 kernel, to model residual red noise in the visit 1 light curves not captured by the aforementioned detrending.
+We train the GP on the timeseries of the second eigenvalue extracted from the PCA, which roughly corresponds to sub-pixel drifts of the trace position on the detector, as we find this reduces the residual scatter in the light curves by a further 10% compared to training the GP on the light curve time axis.
+The characteristic amplitude of the GP was fit independently for each order, but the timescale was shared between both (e.g., Radica et al., 2024).
 During the pre-transit baseline of the second visit, there was a large flare event on the host star.
-We attempted several methods to correct this flare in the light curves e.g., detrending against the H flux, computed for each integration by measuring the area between the flux in the H line and a local quadratic continuum, shown in the second panels in Figure 1.
+We to correct this flare in the light curves by detrending against the H flux, computed for each integration by measuring the area between the flux in the H line and a local quadratic continuum, shown in the second panels in Figure 1.
 However, this avenue was unsuccessful as there is a slight time delay between the appearance of the flare in the H time series and in the white light curve, likely due to lags in energy release as a function of wavelength (Howard et al., 2023).
-We eventually decided that, since the flare occurred in the pre-transit baseline and was sufficiently separated in time from the transit itself, the optimal course of action was to simply cut the affected integration from the fit.
-For visit 2, we, therefore, cut the first 75 integrations of the TSO (denoted with grey shading in Figure 1).
+However, given the broad structure of this flare compared to others seen in TRAPPIST-1 light curves (e.g., Howard et al., 2023), we found that we were able to adequately include the flare in our light curve fits using a GP which again employed the Matérn 3/2 kernel.
+As above, we fit the characteristic amplitude separately to each order, but shared the timescale between both.
+Finally, we include an error inflation term for each order of each visit, which is added in quadrature to the extracted flux errors.
+Our final models consist of 20 parameters for each visit — we use wide and uninformative priors in all cases.
+The results of the white light curve fits are shown in Figure 1 and the best-fitting astrophysical parameters are summarized in Table 1.
 2.3 Spectroscopic Light Curve Fitting
 To construct the spectrophotometric light curves, we sum the first order flux in bins of 80 pixels (corresponding to a constant resolving power of approximately =25).
-The combined effects of the spectrum of the late-M host star, and the low throughput of the SOSS order 2 mean that the second order trace is incredibly faint for observations of TRAPPIST-1 planets (e.g., Lim et al., 2023).
+The combined effects of the spectrum of the late-M host star, and the low throughput of the SOSS order 2 mean that the second order trace is very faint for observations of TRAPPIST-1 planets (e.g., Lim et al., 2023).
 We therefore do not bin order 2 further, and simply include the white light curve results in our transit spectrum.
-The orbital parameters from both visits are consistent between visits (Table 1), and we therefore fix them to the weighted average between the two when fitting the spectroscopically binned light curves.
+The orbital parameters from both visits are consistent between visits (Table 1) as well as with literature values (e.g., Agol et al., 2021), and so we fix them to the weighted average between the two visits when fitting the spectroscopically binned light curves.
 We include the same systematics model as mentioned above for each visit, except that we fix the characteristic timescale of the GP to the best-fitting value from the white light fit for each visit, and allow the amplitude to freely vary, since we can reasonably expect the impacts of stellar variability to be more prominent at bluer vs.
 redder wavelengths (e.g., Radica et al., 2024).
-We find a slight offset of 122 ppm between the spectra of of the two visits, which we attribute to residual dilution from the pre-transit flare in visit 2.
-Indeed, this can also be noticed in the white light curve depths for each visit shown in Table 1, where the visit 2 order 1 transit depth is 100 ppm shallower than visit 1 — suggestive of dilution.
-We tested trimming a different number of pre-transit integrations and found that this slightly changed the offset (to the tune of 5 – 10 ppm), but did not change the shape of the final transit spectrum.
 We also reduce and fit the TSOs from each visit using the independent NAMELESS pipeline (Feinstein et al., 2023; Coulombe et al., 2023).
 Further details of this analysis are presented in Appendix A, and a comparison of the spectra from both pipelines are shown in Figure 7.
-Since the spectra from each pipeline are consistent to within 1- for both visits, we only consider the exoTEDRF spectrum for the remainder of our analysis.
-3 Modelling of the transit light source effect
+Since the spectra from each pipeline are consistent to within 1- for both visits (or 1 via a -per-data-point metric), we only consider the exoTEDRF spectrum for the remainder of our analysis.
+3 Modelling of the Stellar Photosphere
+Before undertaking any planet atmosphere analysis, we first seek to better constrain the properties of TRAPPIST-1’s photosphere, and thereby aid in our attempts to disentangle veritable atmosphere signals from the TLSE.
+To this end, we perform two independent, but complementary analyses: one of the in-transit data via the TLSE (Section 3.1), and another focusing on the out-of-transit data via modelling the stellar spectrum itself (Section 3.2).
+3.1 Transit Light Source Effect Modelling
+. Stellar contamination alone can explain the wavelength-dependent variations in the spectrum, and the retrieved heterogeneity parameters are largely consistent across both visits.
 We first model the transmission spectrum of TRAPPIST-1 c assuming that the TLSE can completely explain each visit’s transmission spectrum.
 This approach is motivated by the presence of small features resembling water absorption (e.g., around 1.4 µm; Figure 2) and a pronounced slope towards short wavelengths in the spectrum of the first visit.
-We use the open-source code stctm (Piaulet-Ghorayeb, 2024) to perform the TLS retrievals.
+We use the open-source code stctm (Piaulet-Ghorayeb, 2024; Piaulet-Ghorayeb et al., 2024) to perform the TLS retrievals.
 We model the presence of two populations of heterogeneities: spots (covering a fraction of the photosphere, and cooler than the photosphere at a temperature ), and faculae (with representing their covering fraction, and their temperature).
@@
 In the retrieval, we select at each iteration the closest-matching model (in effective temperature and surface gravity space) for the photosphere, spot, and faculae components from a finely-sampled pre-computed grid obtained with the MSG module (Townsend & Lopez, 2023) using the PHOENIX (Husser et al., 2013) stellar model grid.
-We also fit the used for the heterogeneities, allowing it to be lower than that of the photosphere following recent work (Lim et al., 2023; Fournier-Tondreau et al., 2024), but retrieve a value consistent with that of the photosphere (Agol et al., 2021).
-We use the affine-invariant Markov Chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013) to explore the parameter space, with the number of walkers set as 20 times the number of fitted parameters, and run the chains for 5000 steps (60% are discarded as burn-in).
+We also introduce the fitting of (Piaulet-Ghorayeb et al., in prep), the difference between the photosphere and heterogeneity components, allowing the heterogeneity spectra to use a lower than that of the photosphere, motivated by recent work (e.g., Fournier-Tondreau et al., 2024).
+We use the affine-invariant Markov Chain Monte Carlo sampler emcee (Foreman-Mackey et al., 2013) to explore the parameter space, with the number of walkers set as 20 times the number of fitted parameters, and run the chains for 5000 steps, visually monitoring for convergence, and discard the first 60% as burn-in.
 In addition to the stellar heterogeneity properties, we fit the wavelength-independent transit depth of TRAPPIST-1 c, .
 We perform independent fits to the spectrum of each visit, as they are separated by about one year and should not, a priori, be affected by the same TLS contribution.
-We find that similar spot and facula properties can explain both visits (Figure 2).
+We also perform additional retrievals considering only faculae or only spots in addition to the quiet stellar photosphere.
+We find that when either only spots or both spots and faculae are included in our model, we obtain a very good match to the observed spectra, without the need for a planetary atmosphere.
+The addition of faculae enables us to fit better the shortest-wavelength data points with a tentative downward slope which cannot be created by spots alone (Figure 2) and results in broader distributions on the spot properties given the interplay with a facula component.
+However, when only faculae are included in our retrievals, the resulting TLS model does not capture the observed spectral features, particularly the broad absorption features around 1.4 and 1.9 µm (as faculae would predict these features to instead be inverted).
+The resulting parameters from these retrievals are therefore not presented, as they result in poorer fits.
+We find that similar spot and facula properties can explain both visits (Figure 2; Table 2), which motivates the use of a shared stellar contamination component in our joint retrievals of stellar and planetary atmosphere contributions to the spectra in Section 4.
+There is, however, a slight () offset between the “bare-rock” transit depths retrieved in the fit to visit 1 ( ppm) and visit 2 ( ppm).
+Finally, our retrievals do not support the use of different values for the photosphere and heterogeneity components, as we only find lower limits on (e.g., Figure 8).
+We use these TLS-only retrievals to produce “TLS-corrected” spectra which we use to explore the limiting effect of stellar contamination on the constraining power of our observations (see Section 5.5; Figure 9), and for visualization purposes in order to interpret the results from our atmosphere retrievals (Figures 4 and 5).
+For each visit, the TLS-corrected spectrum is produced by dividing the observed spectrum by the best-fitting stellar contamination model, and then multiplying the result by the best-fit model bare-rock transit depth (6776 ppm for visit 1, 6730 ppm for visit 2).
+These spectra aim to quantify the precision we could reach on atmospheric inferences given our observing setup if stellar contamination was not a factor.
+3.2 Out-of-Transit Stellar Spectrum Modelling
+In order to verify the reliability of our TLS retrievals described above, as well as gain additional insights into the heterogeneous nature of the photosphere of TRAPPIST-1, we directly model the out-of-transit stellar spectrum, roughly following the methodologies of Wakeford et al.
+(2019) and Moran et al.
+(2023).
+For each visit, we first create a median stack of the stellar spectrum considering only the out-of-transit integrations.
+For visit 1, this includes both the pre- and post-transit baseline.
+For visit 2, we only include integrations before the flare event in the pre-transit baseline.
+The post-transit baseline appears, by eye, to be at a slightly higher level than the pre-transit baseline.
+We test fits for visit 2 using only the pre-transit and only the post-transit baselines, as well as both and note that we do not find any difference in the inferred photosphere or heterogeneity parameters as a result of these choices, and we therefore proceed using both the pre- and pos-transit baseline for the fits.
+We then flux calibrate the stellar spectra, following the procedure presented in Lim et al.
+(2023).
+Like Moran et al.
+(2023), we use the standard deviation of the stellar spectra as a function of time as the errors on each wavelength point, as we find this more representative than the pipeline-extracted flux errors.
+As with the TLS fits, we proceed to fit one-, two-, and three-component models (i.e., photosphere-only, photosphere + spots/facule, photosphere + spots + faculae) directly to the stellar spectra using the StellarFit222https://github.com/radicamc/StellarFit package.
+For consistency with the in-transit fits, we explore a grid of PHOENIX model spectra (Husser et al., 2013), but also test a grid of SPHINX models (Iyer et al., 2023), which include temperatures as low as 2000 K, more appropriate for modelling cold spots on ultra-cool dwarfs like TRAPPIST-1.
+All models in a given grid are scaled by to represent the flux received at Earth, using =0.1192 for the stellar radius (Agol et al., 2021) and =12.1 pc for the distance of the system to the Earth (Gillon et al., 2016).
+One-component fits have four free parameters: the stellar effective temperature, , the surface gravity, as well as a spectrum scaling factor (e.g., Wakeford et al., 2019; Moran et al., 2023), and a multiplicative error inflation term.
+The photosphere temperature was allowed to vary from 2300 – 5000 K for the PHOENIX grid or 2000 – 4000 K for the SPHINX grid, the log gravity from 3.5 – 5.5 (PHOENIX) or 4.0 – 5.5 (SPHINX), and the scale factor from 0.8 – 1.2.
+Multi-component fits also include the spot/faculae temperature (/), covering fraction (/), and the gravity of the heterogeneity (e.g., Lim et al., 2023; Fournier-Tondreau et al., 2024).
+Spot/faculae temperatures are required to be at least 100 K and up to 1000 K cooler/warmer than the photosphere, whereas their gravity must be within 1.0 of the photosphere value.
+Unlike previous works, we do not simply consider a “best-fitting” model, but instead explore the full parameter space of model fits to obtain distributions on the photosphere and heterogeneity parameters analogous to our TLS retrievals.
+We use nested sampling for the posterior exploration, as implemented in the dynesty (Speagle, 2020) package, which also allows us to assess the statistical evidence for including heterogeneities in our models.
+The best fitting three-component models for both stellar grids are shown in Figure 3, and are in excellent agreement with the in-transit TLS results.
+The best-fitting parameters from both the in-transit and out-of-transit analyses are also presented in Table 2.
 4 Joint modelling of atmosphere and stellar contamination
@@
 In the retrieval, the planetary radius is optimized for each set of sampled parameters to obtain the best match to the observed spectrum.
-We perform three categories of retrievals corresponding to different assumptions about the planetary atmosphere: (1) H2/He-dominated scenarios; (2) single-component high-mean molecular weight atmospheres; (3) two-component high-mean molecular weight atmospheres with one infrared absorber present in the atmosphere and either N2 or O2 acting as the background gas.
-We parameterize H2/He-dominated atmospheres in the retrieval with a H2/He mixture as the background gas (assuming He/H2= 0.157), log-uniform priors on the abundances of N2, CH4, H2O, CO, CO2, NH3, and SO2, and a log-uniform prior on the pressure at which the atmosphere becomes opaque (either due to the presence of a cloud deck or a solid surface).
-For single-component high-mean molecular weight atmospheres, we only explore absorbers with prominent features over the NIRISS/SOSS wavelength range; namely, atmospheres made of 100% CH4, H2O, CO, or NH3.
+We perform three categories of retrievals corresponding to different assumptions about the planetary atmosphere: (1) H2/He-dominated scenarios; (2) single-component high-MMW atmospheres; (3) two-component high-MMW atmospheres with one infrared absorber present in the atmosphere and either N2 or O2 acting as the background gas.
+We parameterize H2/He-dominated atmospheres in the retrieval with a H2/He mixture as the background gas (assuming He/H2= 0.157), log-uniform priors on the abundances of N2, CH4, H2O, CO, CO2, and NH3, and a log-uniform prior on the pressure at which the atmosphere becomes opaque (either due to the presence of a cloud deck or a solid surface).
+We use the HELIOS-K (Grimm & Heng, 2015; Grimm et al., 2021) compiled line lists for CH4 (Hargreaves et al., 2020), H2O (Polyansky et al., 2018), CO (Hargreaves et al., 2019), CO2 (Yurchenko et al., 2020), and NH3 (Coles et al., 2019).
+For single-component high-MMW atmospheres, we only explore absorbers with prominent features over the NIRISS/SOSS wavelength range; namely, atmospheres made of 100% CH4, H2O, CO, or NH3.
 We also fit for the “effective surface pressure” at which the atmosphere becomes opaque.
-Finally, we also explore potential N2-dominated atmospheres (similar to some terrestrial objects in the solar system), or O2-dominated atmospheres (motivated by the possible O2-CO2 atmosphere consistent with the MIRI eclipse observations of TRAPPIST-1 c; (Zieba et al., 2023)) where only one of CH4, H2O, CO, or NH3 acts as an absorber, beyond the broad N2-N2 or O2-O2 collision-induced absorption features introduced by the background gas.
+Finally, we also explore potential N2-dominated atmospheres (similar to some terrestrial objects in the solar system), or O2-dominated atmospheres (motivated by the possible O2-CO2 atmosphere consistent with the MIRI eclipse observations of TRAPPIST-1 c; Zieba et al., 2023) where only one of CH4, H2O, CO, or NH3 acts as an additional absorber beyond the broad N2-N2 or O2-O2 collision-induced absorption features.
 In this retrieval setup, we fit for the partial pressure of each component.
-We assume that the vertical temperature structure is isothermal, motivated by the likely-thin region of the upper atmosphere probed by our observations in transmission, and only fit for this photosphere temperature.
-We use the nested sampling method (Skilling, 2004, 2006) to sample the full parameter space, with the multi-ellipsoid method implemented using the nestle module 222https://github.com/kbarbary/nestle within the SCARLET retrieval framework.
+Unfortunately, the CH4 line lists does not extend all the way to the short-wavelength end of NIRISS/SOSS order 2 (Hargreaves et al., 2020).
+Therefore, we exclude the order 2 data points from the spectrum for the 100% CH4, N2-CH4, and O2-CH4 retrievals.
+We assume that the vertical temperature structure is isothermal, motivated by the fact that our transmission spectra would not the precision to measure any temperature gradient in the thin transit photosphere region we are probing.
+We use the nested sampling method (Skilling, 2004, 2006) to sample the full parameter space, with the multi-ellipsoid method implemented using the nestle module333https://github.com/kbarbary/nestle within the SCARLET retrieval framework.
 The final posterior distributions are obtained from retrievals performed using at least 1000 live points.
-We compute forward models at R=15,625 before convolving them to the resolution of the data assuming a uniform throughput within each bin for the likelihood calculation.
-We also calculate forward models for 1 and 100 solar metallicity, H2/He-dominated atmospheres with varying cloud-top pressures (Figure 3; left panel), and models for several high mean molecular weight compositions (Figure 4; top panel), to put our retrieval results in context.
+We compute forward models at R=31,250 before convolving them to the resolution of the spectrum, assuming a uniform throughput within each bin for the likelihood calculation.
+We also calculate forward models for 1 and 100 solar metallicity, H2/He-dominated atmospheres with varying cloud-top pressures (Figure 4; left panel), and models for several high MMW compositions (Figure 5; top panel), to put our retrieval results in context.
 For each model, the temperature-pressure profile is assumed to be isothermal at the equilibrium temperature of TRAPPIST-1 c.
-For each set of retrievals, we account in the modelling for the fact that the transmission spectra we observe are likely heavily affected by stellar contamination (see Figure 2), thereby marginalizing over the interplay that unocculted stellar heterogeneities can have with planetary absorption features.
-In addition to the planetary atmosphere properties, we thus account for stellar heterogeneities assuming a single population of spots, fitting for the spot covering fraction and their temperature contrast with the photosphere, as well as the photosphere temperature itself.
-We also assume that the stellar heterogeneity components of both visits are identical, motivated by the results from the stctm fit (Figure 2), and do not fit for a different for the heterogeneity component.
-We note, though, that we performed one set of atmosphere+stellar contamination retrievals where we fit for visit-specific stellar heterogeneity properties, and found that this does not affect our inferences, further motivating our choice of shared spot parameters for both visits.
-We place uniform priors on the spot fraction and temperature, but use the same Gaussian prior as in the TLS-only fit for the photosphere temperature.
-We note that we find a vertical offset of 100 ppm between the spectra of both visits (as can be inferred from the white light transit radii presented in Table 1).
-We attribute it to the presence of the flare in the pre-transit baseline of the second visit that may result in a dilution effect when fitting the transit light curves and therefore fit for a parameter describing the visit 1-visit 2 offset.
+For each set of retrievals we account for the impacts of stellar contamination, thereby marginalizing over the interplay between unocculted stellar heterogeneities and planetary absorption features.
+In addition to the planetary atmosphere properties, Motivated by the results in Section 3, we assume two heterogeneity populations: spots (cooler than the photosphere), and faculae (hotter than the photosphere).
+For each heterogeneity component, we fit a covering fraction (with a wide uniform prior of 0 to 50%) and a temperature contrast to the photosphere (following the uniform priors of the TLS-only fit).
+We also fit for the photosphere temperature itself, with a Gaussian prior based on the value reported by Agol et al.
+(2021).
+We allow the photosphere to vary within the range as lower values are required to obtain a good match of the PHOENIX models to the out-of-transit spectrum of TRAPPIST-1 (see e.g., Lim et al., 2023 and Section 3.2).
+Given that the TLS-only fit did not provide evidence for a difference in between the photosphere and heterogeneity components, we fix them to the same value for each model evaluation.
+Finally, we fit for an offset between the visit 1 and visit 2 spectra, motivated by the stctm retrieval results.
+Given the consistency between the stellar heterogeneity parameters inferred from both the TLS and stellar spectrum fits, we assume that the heterogeneity components for both visits are identical in our joint atmosphere fits.
+However, we note that we also performed one set of atmosphere+stellar contamination retrievals where we fit for visit-specific stellar heterogeneity properties, and found that this does not affect our inferences.
+We show the results for H2/He-dominated atmospheres in Figure 4, and high-MMW atmospheres in Figures 5 and 6.
 5 Discussion
-5.1 Impacts of Stellar Contamination
-Our two-transit study of TRAPPIST-1 c aptly summarizes the promise and peril of transmission observations of rocky planets around M dwarf host stars.
-There are now a plethora of studies that have shown that stellar contamination is a true bottleneck in the attempt to detect atmospheres around rocky, M dwarf planets via transmission spectroscopy (e.g., May et al., 2023; Moran et al., 2023; Lim et al., 2023; Cadieux et al., 2024).
-In most cases, different manifestations of the TLSE (i.e., different temperature distributions, covering fractions, etc.
-of stellar spots and faculae) give rise to obviously different transmission spectra between visits, allowing for a clear inference of the impacts of the TLSE, even without undertaking in-depth modelling.
-Consistency in the location and shapes of spectral features between visits has been argued to lend credence to an interpretation favouring a planetary atmosphere origin (e.g.
-May et al., 2023).
-However, despite the high level of consistency between the transmission spectra (1.1- on average), our retrieval analysis still finds that both visits can be entirely explained by stellar contamination.
-As shown in Figure 2, we find very similar stellar heterogeneity parameters (spot/faculae temperatures and covering fractions) for both visits.
-The spot and faculae temperatures are consistent with those found by Lim et al.
-(2023) in their analysis of NIRISS/SOSS transmission spectra of TRAPPIST-1 b.
-However, we find both visits to be spot-dominated, with minimal contributions from faculae, whereas Lim et al.
-(2023) have one spot-dominated, and one facula-dominated visit.
-The consistency between visits would perhaps be unsurprising if they were separated by a very small period of time, which would therefore probe similar spot and facula distributions on the host star.
-However, our two visits are separated by 367 d, significantly longer than the 3.3 d stellar rotation period (Luger et al., 2017).
-When phase-folded to the stellar rotation period, this corresponds to a phase difference of 0.21 — similar to the observations of Lim et al.
-(2023), who, as mentioned above, found no such consistency between their two visits.
-This highlights the fact that TLSE models are generally non-predictive of future stellar contamination, since the distributions of spots and faculae in the host star’s photosphere is entirely stochastic (Rackham et al., 2018, 2023; TRAPPIST-1 JWST Community Initiative et al., 2024).
+5.1 Inferences of the Stellar Photosphere
+For the out-of-transit stellar modelling using both grids and visits, the photosphere temperature is generally higher than, but still broadly consistent with literature values (e.g., Agol et al., 2021; Davoudi et al., 2024).
+The retrieved log gravities (3.8 for PHOENIX and 4.2 for SPHINX) are both significantly lower than the 5.2 found by Agol et al.
+(2021) and Davoudi et al.
+(2024), through this discrepancy was also noted in the TRAPPIST-1 spectra obtained with NIRISS/SOSS by Lim et al.
+(2023) in their observations of planet b.
+Our results from both grids are broadly consistent with the TLS retrievals (e.g., Figures 2, 3; Table 2).
+For SPHINX, a two-component model including a photosphere and spots is preferred at 4 (using the Benneke & Seager (2013) Bayesian evidence to -level conversion) over a homogeneous photosphere model and at 2 over a model including both spots and faculae for both visits.
+Like in the TLS retrievals, when faculae are included, the fits push the covering fraction to zero in order to minimize their contribution.
+Both the spot covering fraction and temperature are in excellent agreement with the TLS results.
+The spot temperatures are also consistent with those inferred by previous studies (e.g., Wakeford et al., 2019; Garcia et al., 2022; Lim et al., 2023; Davoudi et al., 2024).
+For the PHOENIX grid, the results are also in excellent agreement with both the SPHINX grid and the TLS fit results (Table 2).
+In this case, the three-component fit is preferred (though the significance is low; 1.8).
+The faculae covering fraction is also only 3%, and therefore essentially negligible.
+For all grids tested, we find that the gravity of the heterogeneities is always consistent with that of the photosphere.
+We note though that the overall quality of the fits are unsatisfactory, with values generally 350.
+The short wavelengths (1.1 µm) are especially badly fit by the PHOENIX grid in particular.
+In fact, the PHOENIX grid struggled to fit both the short and long wavelengths simultaneously, routinely overestimating the flux 1.7 µm and finding photosphere temperatures that were highly inconsistent with literature values.
+We therefore restricted the models to wavelengths 1.1 µm for the PHOENIX grid where the models were better matches to the data.
+Performing the same test for the SPHINX grid does not alter the retrieved parameter distributions.
+Another caveat is that the high resolution of NIRSpec/G395H (2700) allowed Moran et al.
+(2023) to fit to individual stellar lines.
+However, in our case, the medium resolution of NIRISS/SOSS (700) does not provide this same capability (see the inset in Figure 3), and our fits are more sensitive to the broadband structures in the spectrum.
+In all, though we confirm the results of the in-transit TLS fits, our stellar spectrum analysis demonstrates the clear need for improved stellar models for ultra-cool stars.
+Moreover, we posit that since the in-transit analysis relies on ratios of stellar models instead of absolute fits, it may be more robust to the aforementioned model inaccuracies than direct fits to stellar spectra — something that should be explored further.
 5.2 Cloud-Free and Thick Hydrogen-Rich Atmospheres Are Ruled Out
-We find that, similarly to the closer-in planet TRAPPIST-1 b (Lim et al., 2023), we can confidently rule out thick or cloud-free hydrogen-dominated atmospheres, with a 3- upper limit of only 0.1 bar on the effective surface pressure of a H2/He-dominated atmosphere atop TRAPPIST-1 c (Figure 3).
-The thin 0.1 bar hydrogen-rich atmosphere scenarios allowed by the transmission spectrum are, moreover, unlikely from an atmospheric evolution standpoint — TRAPPIST-1 c was exposed to large cumulative amounts of high-energy XUV irradiation from its host star, sufficient to strip away any primordial atmosphere (Hori & Ogihara, 2020; Turbet et al., 2020) even when considering hydrogen outgassing (Hu et al., 2023).
-The transmission spectrum still allows within the 3- confidence contours for thick, high-metallicity (100 solar) atmospheres, which may be more resilient against escape due to their higher mean molecular weights.
-5.3 Limits on High Mean Molecular Weight Atmospheres
-We, therefore, turn to evaluate the range of plausible high-metallicity atmosphere scenarios for TRAPPIST-1 c, and focus on the detectability of H2O-, NH3-, CO- and CH4-rich atmospheres, as these infrared absorbers have sufficient opacity within the NIRISS/SOSS wavelength range to obtain meaningful constraints.
-We can rule out thick (1 bar) 100% H2O and NH3 atmospheres at better than 3, and even a 1 bar 100% CO atmosphere is excluded at the 2- level, although thinner 0.1 bar atmospheres remain plausible explanations for the flat spectrum (Figure 4).
-With the exception of CH4, we find tight upper limits on the abundance of all the other molecules tested (H2O, NH3, CO), even in lower-mean molecular weight atmospheres dominated by N2 or O2 as the background gas (Figure 5).
-We compare those limits, where available, with measured compositions of solar system planets.
+We find that, similarly to the closer-in planet TRAPPIST-1 b (Lim et al., 2023), we can confidently rule out thick or cloud-free hydrogen-dominated atmospheres, with a 3- limit of 10 mbar on the effective surface pressure of a H2/He-dominated atmosphere atop TRAPPIST-1 c (Figure 4).
+The thin hydrogen-rich atmosphere scenarios allowed by the transmission spectrum are, moreover, unlikely from an atmospheric evolution standpoint — TRAPPIST-1 c was exposed to large cumulative amounts of high-energy XUV irradiation from its host star, sufficient to strip away any primordial atmosphere (Hori & Ogihara, 2020; Turbet et al., 2020) even when considering hydrogen outgassing (Hu et al., 2023).
+The transmission spectrum still allows within the 3- confidence contours for thick, high-metallicity (100 solar) atmospheres, which may be more resilient against escape due to their higher MMWs.
+5.3 Limits on High-Mean Molecular Weight Atmospheres
+We, therefore, turn to evaluate the range of plausible high-metallicity atmosphere scenarios for TRAPPIST-1 c, and focus on the detectability of H2O-, NH3-, CO- and CH4 in high-MMW atmospheres, as these infrared absorbers have sufficient opacity within the NIRISS/SOSS wavelength range to enable meaningful constraints.
+As expected from the high-metallicity tail of the surface pressure-metallicity distribution (Figure 4), the densest atmosphere cases we tested do not provide strong upper limits on the effective surface pressure.
+We find that deep, 100% H2O, NH3, and CH4 atmospheres remain compatible with our spectrum at the 3 level, although 1 bar atmospheres are disfavored at the 2 level (Figure 5).
+In the 100% CO case, we exclude even a 1-bar atmosphere at the 3 level.
+However, we caution that this result mostly stems from the fact that CO has minimal opacity in the SOSS waveband from 2–2.5 µm, and a pure-CO atmosphere would elsewhere yield a perfectly flat spectrum.
+Therefore, this example only rules out 100% CO atmospheres, but likely not cases where a CO-rich atmosphere is mixed with small amounts of another infrared absorber.
+We then turn to models that allow for lower atmospheric MMWs, where the infrared absorber can be a trace species rather than a dominant component — representative of the atmospheres of some solar system planets.
+In this scenario, the background gas is N2 or O2, with only one of the aforementioned species as the active infrared absorber (Figure 6).
+We compare these limits, where available, with measured compositions of solar system bodies.
 In particular, we highlight in Figure 6 the amount of H2O in Earth’s N2-dominated atmosphere (using 0.4% as the mean mole fraction over the entire 1-bar atmosphere, although it can reach about one percent at sea level; Kasai et al., 2011; Kelsey et al., 2022), or the amount of CH4 in Titan’s atmosphere, assuming either a cloud-free Titan scenario (where one can see down to the surface) with 5% CH4, or a hazy scenario accounting for the fact that hydrocarbon haze opacity raises the effective surface pressure to 10 mbar, and with a factor of three less methane in the upper atmosphere (see e.g.
-Charnay et al., 2014.
-We find that similar amounts of water to what is observed in Earth’s atmosphere are ruled out for TRAPPIST-1 c, while the transmission spectrum alone is not sensitive enough to the CH4 abundance to offer meaningful constraints on Titan-like atmospheres.
-Although our observations cannot rule out the presence of CH4 either in thick 1–100 bar pure-CH4 atmospheres, or as a trace species in a N2- or O2-dominated atmosphere, CH4 is unlikely on TRAPPIST-1 c from a modelling standpoint because of its vulnerability to destruction via photochemistry (Turbet et al., 2018).
-Since O2 is a heavier molecule than N2, the O2 background scenarios generally correspond to more compact, higher mean molecular weight atmospheres.
+Charnay et al., 2014).
+We find that similar amounts of water to what is observed in Earth’s atmosphere are disfavored at the 2 level for TRAPPIST-1 c, while the transmission spectrum alone is not sensitive enough to the CH4 abundance to offer meaningful constraints on Titan-like atmospheres.
+Although our observations cannot rule out atmospheres where NH3 or CH4 make up the entire volatile budget (or are trace species in N2- or O2-dominated atmospheres), the presence of methane and ammonia is unlikely on TRAPPIST-1 c from a modelling standpoint because of its vulnerability to destruction via photochemistry (Turbet et al., 2018).
+Since O2 is a heavier molecule than N2, the O2 background scenarios generally correspond to more compact, higher-MMW atmospheres.
 Therefore, although the results are qualitatively similar to the N2-background constraints, our transmission spectrum allows for larger amounts of the tested molecular absorbers in an O2-dominated atmosphere.
-As such, we unfortunately are not able to rule on any of the possible scenarios presented by Zieba et al.
+Further, because of the lack of sensitivity to CO2 absorption with NIRISS/SOSS spectroscopy, we are not able to rule on any of the possible scenarios presented by Zieba et al.
 (2023) based on their MIRI eclipse observations of this planet.
@@
 5.4 Implications for the Presence of an Atmosphere on TRAPPIST-1c
-Ultimately, our results from the analysis of the observed transmission spectrum of TRAPPIST-1 c are in line with modeling predictions that it would not have been able to retain any substantial atmosphere over the 5–9 Gyr lifetime of its host star (Burgasser & Mamajek, 2017), regardless of its initial volatile content (Krissansen-Totton, 2023).
-The thin, 1 bar atmospheres that remain allowed by our observations, even if they correspond to high mean molecular weights, are unlikely to be retained on TRAPPIST-1 c from a theoretical perspective.
-Even if Jeans escape (generally less efficient than blow-off) is expected to dominate for high-mean molecular weight atmospheres atop TRAPPIST-1 c, the escape rates even for resilient N2-CO2 atmospheres range from one Earth atmosphere per Myr to one Earth atmosphere mass per Gyr (Van Looveren et al., 2024) at the present-day irradiation of this planet.
+Although it could still be explained by some high metallicity scenarios, our non-detection of any significant atmospheric signature in the transmission spectrum of TRAPPIST-1 c is also in line with modeling predictions that TRAPPIST-1 c could be a bare rock, as any substantial atmosphere is expected to be stripped over the 5–9 Gyr lifetime of its host star (Burgasser & Mamajek, 2017), regardless of its initial volatile content (Krissansen-Totton, 2023).
+Even if they remain tentative (at the 2 level), our atmospheric retrieval results disfavor thick 1 bar atmospheres, even for high MMW compositions.
+The thin atmosphere scenarios that remain consistent with our observations given their smaller spectral features are however unlikely to be retained on TRAPPIST-1 c from a theoretical perspective.
+Even if Jeans escape (generally less efficient than blow-off) is expected to dominate for high-MMW atmospheres atop TRAPPIST-1 c, the escape rates even for resilient N2-CO2 atmospheres range from one Earth atmosphere per Myr to one Earth atmosphere mass per Gyr (Van Looveren et al., 2024) at the present-day irradiation of this planet.
 The implications are even stricter when considering the fact that any atmosphere on TRAPPIST-1 c also would have had to withstand the early XUV-active stages of the late M dwarf’s evolution (Bolmont et al., 2017), and its position inside of the runaway greenhouse limit (Turbet et al., 2020) indicates that water could not have been efficiently shielded from escape via condensation in a liquid water ocean, even if this limit is pushed to closer host star-planet distances around M dwarfs in 3D models (Turbet et al., 2023).
+5.5 Impacts of Stellar Contamination on the Transmission Spectra
+Our two-transit study of TRAPPIST-1 c aptly summarizes the promise and peril of transmission observations of rocky planets around M dwarf host stars.
+There are now a plethora of studies that have shown that stellar contamination is a true bottleneck in the attempt to detect atmospheres around rocky, M dwarf planets via transmission spectroscopy (e.g., Zhang et al., 2018; Wakeford et al., 2019; May et al., 2023; Moran et al., 2023; Lim et al., 2023; Cadieux et al., 2024; Rackham & de Wit, 2024).
+In most cases, different manifestations of the TLSE (i.e., different temperature distributions, covering fractions, etc.
+of stellar spots and faculae) give rise to obviously different transmission spectra between visits, allowing for a clear inference of the impacts of the TLSE, even without undertaking in-depth modelling.
+Consistency in the location and shapes of spectral features between visits has been argued to lend credence to an interpretation favouring a planetary atmosphere origin (e.g.
+May et al., 2023).
+However, despite the high level of consistency between the transmission spectra (1.1- on average), our retrieval analysis still finds that both visits can be entirely explained by stellar contamination.
+As shown in Figure 8, and summarized in Table 2, we find very similar stellar heterogeneity parameters (spot/faculae temperatures and covering fractions) for both visits.
+We also find both visits to be spot-dominated, with minimal contributions from faculae, whereas Lim et al.
+(2023) have one spot-dominated, and one facula-dominated visit in their two-transit analysis of TRAPPIST-1 b.
+The consistency between visits would perhaps be unsurprising if they were separated by a very small period of time, which would therefore probe similar spot and facula distributions on the host star.
+However, our two visits are separated by 367 d, significantly longer than the 3.3 d stellar rotation period (Luger et al., 2017).
+When phase-folded to the stellar rotation period, this corresponds to a phase difference of 0.21 — similar to the observations of Lim et al.
+(2023), who, as mentioned above, found no such consistency between their two visits.
+Moreover, both of our visits are at comparable stellar phases to those of Lim et al.
+(2023).
+This highlights the fact that the distributions of spots and faculae in the host star’s photosphere are entirely stochastic, and therefore TLSE models are non-predictive of future stellar contamination.
+Irrespective of whether we perform TLS-only or joint TLS-atmosphere retrievals, we still infer consistent stellar contamination parameters (Figure 8).
+This is consistent with our finding that the retrieval does not pick up on any “true” atmosphere signature that would warrant changes in the inferred TLS parameters.
+Finally, we evaluated the limiting effect of having to marginalize over stellar contamination on our sensitivity to thin 100% H2O atmospheres.
+This choice was motivated by the fact that water is the molecule most affected by this marginalization, given that its features coincide with prominent unocculted stellar spot signatures, but refer the reader to Piaulet-Ghorayeb et al.
+(in prep) for a more detailed exploration of the impact of TLSE marginalization on other infrared absorbers.
+We performed a retrieval on the TLS-corrected spectra from both visits (see Section 3.1), using the exact same setup as the joint atmosphere + TLS retrieval except that only the atmosphere properties are retrieved.
+We find that our sensitivity drops by about two orders of magnitude through accounting for stellar contamination.
+As shown in Figure 9, without stellar contamination we could rule out at the 2 (3) level 100% H2O atmospheres with surface pressures down to 1 mbar (1.6 bar).
+However, when marginalizing over the TLSE, we can only rule out at the 2 (3) level cases with surface pressures of less than 0.1 bar (45 bar).
+Therefore, we are in a regime where the unknown TLS contribution to the spectrum, rather than the precision of the measured transit depths themselves, is the limiting factor in our sensitivity to the presence of high-MMW atmospheres.
+Furthermore, this effect will be even more problematic in other datasets where one needs to marginalize over different stellar contamination realizations for each visit in joint retrievals (e.g., Piaulet-Ghorayeb et al., in prep).
 6 Conclusions
@@
 Our 0.6 – 2.85 µm NIRISS/SOSS spectrum reveals strong contamination by the TLSE in both visits.
-However, unlike in previous multi-visit analyses of rocky planets around M dwarfs (e.g., May et al., 2023; Moran et al., 2023; Cadieux et al., 2024), including TRAPPIST-1 b (Lim et al., 2023), we find consistent stellar heterogeneity properties between our two visits, resulting in nearly identical manifestations of the TLSE.
+However, unlike in previous multi-visit analyses of rocky planets around M dwarfs (e.g., May et al., 2023; Moran et al., 2023; Cadieux et al., 2024), including TRAPPIST-1 b (Lim et al., 2023), we find consistent stellar heterogeneity properties between our two visits.
+Via an analysis of the in-transit data (i.e., the TLSE) and direct modelling of the out-of-transit stellar spectrum, we find the stellar photosphere to be dominated by spots 300 K cooler than the ambient photosphere temperature and with a covering fraction of 25%.
+Faculae have negligible contributions to both visits, irrespective of the analysis technique or choice of stellar models.
+This results in nearly identical manifestations of the TLSE and thus transmission spectra.
 Our results support the need for joint modelling of the TLSE and atmosphere properties to achieve robust constraints on rocky, M dwarf planets in the era of JWST.
 Importantly, further stellar model fidelity is critically required to accurately model out the impact of the TLSE and reach the sensitivity to detect or rule out thin atmospheres on small planets via their transit spectra.
-When accounting for stellar contamination in our atmosphere analysis, we find that thick, clear hydrogen-dominated compositions are ruled out at better than 3- significance, and even high-mean molecular weight compositions rich in H2O, NH3 or CO are excluded for an atmosphere thickness of 1 bar or more at better than 2.
-The NIRISS/SOSS transmission spectrum of TRAPPIST-1 c alone cannot rule out all high mean molecular weight compositions, especially for CH4- and CO2-rich cases, to which our spectrum is not highly sensitive.
-However, in light of modelling predictions that CH4-rich compositions are unlikely as CH4 is sensitive to photochemical destruction (Turbet et al., 2018), and considering that Jeans escape would quickly strip TRAPPIST-1 c of even CO2-rich atmospheres (Van Looveren et al., 2024), we conclude that this planet is unlikely to bear any atmosphere whatsoever.
+When accounting for stellar contamination in our atmosphere analysis, we find that thick, clear hydrogen-dominated compositions are ruled out at better than 3- significance.
+The level of precision we reach does not enable us to rule out the cases with the highest MMW, where the atmosphere would be composed of 100% of a heavy volatile species such as H2O, NH3, or CH4, but even these scenarios are disfavored for 1 bar atmospheres at the -level due to the flatness of the TLSE-corrected spectrum.
+The sensitivity increases to partial pressures of about 10 mbar at 2 for the same species in atmospheres with N2 or O2 as the background gas.
 Regardless, as shown by Krissansen-Totton (2023), even if TRAPPIST-1 c is completely atmosphere-less, this does not imply a similar fate for the outer, habitable-zone planets.
@@
 (2024).
-We flag bad pixels by looking for pixels whose second derivative are significant outliers in the spatial domain.
+We flag bad pixels by looking for pixels whose second derivative are significant (7) outliers in the spatial domain.
 The counts of all flagged pixels are then corrected using the bicubic interpolation method of the scipy.interpolate.griddata function.
-We correct for the non-uniform background by individually scaling the two regions of the STScI model background333https://jwst-docs.stsci.edu/ that are separated by the sudden jump in background flux situated around spectral pixel x700.
+We correct for the non-uniform background by individually scaling the two regions of the STScI model background444https://jwst-docs.stsci.edu/ that are separated by the sudden jump in background flux situated around spectral pixel x700.
 Any remaining cosmic rays are corrected by clipping any count that is more than four standard deviations away from the running median for all pixels.
@@
 We then extract the stellar spectra from both spectral orders using a simple box aperture with a width of 36 pixels.
-We perform the broadband and spectroscopic fits of the light curves produced by the NAMELESS reduction with the ExoTEP framework (Benneke et al., 2019a, b).
+We perform the broadband and spectroscopic fits of the light curves with the ExoTEP framework (Benneke et al., 2019a, b).
 For our fit of the white-light curve, which is obtained by summing all wavelengths of the first spectral order (0.85–2.85 m), we keep free the mid-transit time (, considering uniform priors spanning the time of the observations for each visit), planet-to-star radius ratio (, [0.01,0.2]), semi-major axis (, [15,45]), and the impact parameter (, [0,1]).
@@
 The resulting spectra, binned to a fixed resolving power of , are shown in Figure 7.
-Appendix B Additional Tables
+Appendix B Additional Figure & Tables
 Table 1 contains the best-fitting white light curve parameters for both NIRISS/SOSS visits.
 The orbital parameters (, ) are consistent with each other, and with previous literature (e.g., Agol et al., 2021).
-| Parameter | Value | ||
-| Visit 1 | Visit 2 | Weighted Average | |
-| T0 [BJD] | 2459881.401521 | 2460249.515096 | - |
-| 0.133 | 0.136 | ||
-| 28.31 | 28.12 | ||
-Notes: For each visit, and the limb darkening parameters ( and ) were fit separately for each order. All other parameters were jointly fit to both.
+Parameter | Prior Range | Value | ||
+Visit 1 | Visit 2 | Weighted Average | ||
+T0 [BJD] | [T02 hr] | 2459881.401528 | 2460249.515031 | |
+[0.01, 0.9] | 0.08346 | 0.08344 | ||
+[0.01, 0.9] | 0.08364 | 0.08415 | ||
+[70, 90] | 89.73 | 89.38 | ||
+[5, 50] | 28.35 | 28.18 | ||
+[-1, 1] | 0.405 | 0.283 | ||
+[-1, 1] | 0.168 | 0.346 | ||
+[-1, 1] | 0.671 | 0.302 | ||
+[-1, 1] | 0.098 | 0.105 |
+Note: For each visit, and the limb darkening parameters ( and ) were fit separately for each order. All other parameters were jointly fit to both.
+Table 2 contains the best-fitting parameters from the three-component stellar models via both the in-transit (TLS) and out-of-transit analyses.
+Parameter | Model | In-Transit (TLS) | Out-of-Transit | ||
+Visit 1 | Visit 2 | Visit 1 | Visit 2 | ||
+Teff [K] | PHOENIX | ||||
+SPHINX | |||||
+PHOENIX | |||||
+SPHINX | |||||
+[K] | PHOENIX | 2437 | 2432 | ||
+SPHINX | |||||
+PHOENIX | |||||
+SPHINX | |||||
+PHOENIX | |||||
+SPHINX | |||||
+[K] | PHOENIX | ||||
+SPHINX | 3054 | 3133 | |||
+PHOENIX | |||||
+SPHINX | 0.07 | 0.05 | |||
+PHOENIX | |||||
+SPHINX |
+Note: For the stctm fits, the values quoted for and are derived from the posterior samples of , , and , and the intervals reported for the heterogeneity are derived from the and samples.
 Table 3 has the exoTEDRF transmission spectrum used in this work.
@@
 (µm) | (µm) | (ppm) | (ppm) |
-| 0.72 | 0.12 | 6050.99 473.23 | 6982.25 418.63 |
-| 0.90 | 0.04 | 6585.89 111.83 | 6757.96 135.93 |
-| 0.98 | 0.04 | 6772.63 91.97 | 6808.19 128.63 |
-| 1.05 | 0.04 | 6881.03 71.42 | 6960.16 89.30 |
-| 1.12 | 0.04 | 6820.05 69.73 | 6849.88 64.50 |
-| 1.20 | 0.04 | 6699.58 63.55 | 6907.84 80.39 |
-| 1.28 | 0.04 | 6694.65 64.58 | 6697.34 90.68 |
-| 1.35 | 0.04 | 6971.35 60.31 | 7023.65 58.31 |
-| 1.43 | 0.04 | 6886.42 76.10 | 7001.35 76.95 |
-| 1.51 | 0.04 | 6847.72 84.78 | 6937.97 88.64 |
-| 1.59 | 0.04 | 6789.90 76.61 | 6902.63 90.38 |
-| 1.66 | 0.04 | 6895.75 74.07 | 7088.25 92.70 |
-| 1.74 | 0.04 | 6914.99 90.90 | 6899.64 100.01 |
-| 1.82 | 0.04 | 7006.82 97.16 | 7136.91 144.17 |
-| 1.90 | 0.04 | 6804.84 126.15 | 6967.40 122.15 |
-| 1.98 | 0.04 | 7032.96 118.65 | 7127.79 126.21 |
-| 2.06 | 0.04 | 6796.19 132.58 | 6970.70 136.08 |
-| 2.14 | 0.04 | 6842.86 136.57 | 6972.46 134.27 |
-| 2.21 | 0.04 | 7010.63 135.26 | 6977.26 147.22 |
-| 2.29 | 0.04 | 6919.63 128.65 | 6890.04 155.09 |
-| 2.37 | 0.04 | 6707.69 153.51 | 6972.92 152.57 |
-| 2.45 | 0.04 | 6643.73 180.88 | 6821.97 202.91 |
-| 2.53 | 0.04 | 6770.44 232.33 | 7296.84 270.63 |
-| 2.61 | 0.04 | 6837.59 233.79 | 6966.78 293.63 |
-| 2.69 | 0.04 | 6183.53 235.11 | 6961.30 273.61 |
-| 2.77 | 0.04 | 6589.56 270.29 | 6630.95 311.46 |
+0.72 | 0.12 | 6495.76 463.23 | 7080.25 918.63 |
+0.90 | 0.04 | 6940.96 162.46 | 6884.63 193.99 |
+0.98 | 0.04 | 6863.36 112.06 | 7026.99 147.35 |
+1.05 | 0.04 | 7008.37 98.91 | 7049.40 95.48 |
+1.12 | 0.04 | 6999.78 83.35 | 6889.46 77.15 |
+1.20 | 0.04 | 6952.18 71.99 | 6988.60 74.63 |
+1.28 | 0.04 | 6957.36 70.95 | 6955.08 72.35 |
+1.35 | 0.04 | 7205.70 81.75 | 7081.42 78.11 |
+1.43 | 0.04 | 7093.85 98.13 | 7079.19 95.84 |
+1.51 | 0.04 | 7103.09 94.84 | 7154.09 93.00 |
+1.59 | 0.04 | 7066.62 97.78 | 6949.11 120.23 |
+1.66 | 0.04 | 7216.39 96.84 | 7171.97 95.48 |
+1.74 | 0.04 | 7195.63 109.50 | 7011.15 110.87 |
+1.82 | 0.04 | 7206.13 127.63 | 7295.62 123.62 |
+1.90 | 0.04 | 7057.07 145.74 | 7217.11 158.78 |
+1.98 | 0.04 | 7342.25 148.19 | 6961.31 144.52 |
+2.06 | 0.04 | 7004.14 175.93 | 7073.29 167.36 |
+2.14 | 0.04 | 7160.88 151.13 | 7154.43 157.82 |
+2.21 | 0.04 | 7394.33 166.09 | 6897.93 147.24 |
+2.29 | 0.04 | 6970.20 153.40 | 6997.21 172.76 |
+2.37 | 0.04 | 6874.09 223.18 | 6547.71 198.38 |
+2.45 | 0.04 | 7077.64 225.76 | 7116.19 245.70 |
+2.53 | 0.04 | 7046.87 340.02 | 7388.53 327.30 |
+2.61 | 0.04 | 7468.39 389.34 | 7336.15 420.91 |
+2.69 | 0.04 | 6707.51 385.19 | 6884.10 356.58 |
+2.77 | 0.04 | 6621.37 414.79 | 6666.19 417.26 |
+We compare the posterior distributions of the spot, facula, and photosphere properties inferred from the TLS-only fit to the visit 1 and visit 2 spectra, and the joint atmosphere and stellar contamination retrievals (for the single-component atmosphere cases) in Figure 8.
+Finally, we also present the results of our sensitivity study assessing the impact of marginalizing over the TLS effect on the atmosphere thicknesses that can be ruled out by our spectrum, in Figure 9.